Two numbers differ by 3. Their squares differ by 189.
Here is what I have come up with but cannot solve:
(x+3)^2 - x^2 = 189
Am I on the right track?

Answer by ankor@dixie-net.com(12704)

Two numbers differ by 3. Their squares differ by 189.
:
(x+3)^2 - x^2 = 189; is right on!
:
Just FOIL (x+3)(x+3)
x^2 + 6x + 9 - x^2 = 189
:
The neat thing is that the x^2 eliminate themselves, so we have
6x + 9 = 189
6x = 189 - 9
6x = 180
x = 30
;
:
Check this 33^2 - 30^2
1089 - 900 = 189!!